Use the like-bases property and exponents to solve the equation I-8 3 = 324 +7 1 Ox= 3 Ox = - 15 Ox= 1 Ox= -5
Answer #1

2-8 con 2x+7 3 3. on Log(6+)) = log (3**?) both sides ( 3²x+7) || Take log (1-0) of log :)) =(22+7) 10g () 7-8 = (log (32) +